L-1{(e-2s)/[s2(s-1)]}


1
Expert's answer
2021-09-07T02:09:39-0400
"\dfrac{1}{s^2(s-1)}=\dfrac{A}{s}+\dfrac{B}{s^2}+\dfrac{C}{s-1}"

"=\dfrac{As(s-1)+B(s-1)+Cs^2}{s^2(s-1)}"

"=\dfrac{As^2-As+Bs-B+Cs^2}{s^2(s-1)}"

"s^2:A+C=0"

"s^1:-A+B=0"

"s^0:-B=1"

"A=-1, B=-1, C=1"

"L^{-1}(\dfrac{e^{-2s}}{s^2(s-1)})=H(t-2)L^{-1}(-\dfrac{1}{s}-\dfrac{1}{s^2}+\dfrac{1}{s-1})(t-2)"

"=H(t-2)(-H(t-2)-(t-2)+e^{t-2})"



"L^{-1}(\dfrac{e^{-2s}}{s^2(s-1)})=H(t-2)(-H(t-2)-(t-2)+e^{t-2})"